# Aegis: A Decentralized Expansion Blockchain

### v2

@@
Abstract
Blockchains implement monetary systems operated by committees of nodes. The robustness of established blockchains presents an opportunity to leverage their infrastructure for creating expansion chains. Expansion chains can enhance scalability, provide additional functionality to the primary chain they leverage, or implement separate functionalities, while benefiting from the primary chain’s security and the stability of its tokens. Indeed, tools like Ethereum’s EigenLayer enable nodes to stake (deposit collateral) on a primary chain to form a committee responsible for operating an expansion chain.
But here is the rub. Classical protocols assume correct, well-behaved nodes stay correct indefinitely. Yet in our case, the stake incentivizes correctness—it will be slashed (revoked) if its owner deviates. Once a node withdraws its stake, there is no basis to assume its correctness.
@@ Abstract
To address the new challenge, we present Aegis, an expansion chain based on primary-chain stake, assuming a bounded primary-chain write time. Aegis uses references from Aegis blocks to primary blocks to define committees, checkpoints on the primary chain to perpetuate decisions, and resets on the primary chain to establish a new committee if the previous one becomes obsolete. It ensures safety at all times and rapid progress when latency among Aegis nodes is low.
1 Introduction
@@
Blockchains like Ethereum [63] implement monetary systems and allow users to deploy smart contracts [58] with arbitrary logic. The users issue transactions to exchange assets and interact with smart contracts. Blockchain nodes aggregate transactions into blocks, forming a chain that determines the system’s state. However, as blockchain adoption grows, scalability has emerged as a critical challenge [19, 65]. Blockchains have limited throughput, often high latency, and are limited to deterministic execution.
To address these scalability issues and implement new functionality, numerous systems expand a primary blockchain in various ways [37]. Some systems improve scalability by offloading computation, thereby enhancing performance and reducing costs [56, 31]. Other expansion systems serve complementary purposes. These include oracles that aggregate external data for smart contract interactions [18, 26], and mechanisms for generating verifiable randomness [17].
In some cases, it is possible to implement an expansion as a separate Proof of Stake (PoS) blockchain. PoS protocols allow users to stake—lock their tokens as collateral [34, 21, 22]. Until they unstake—withdraw the collateral—they are active nodes, allowed to create new blocks. If they misbehave, they are penalized via slashing [14], losing all or part of their collateral. This approach works well for major blockchains that have reached a critical mass of participants and assets [1]. However, smaller nascent systems do not have sufficient reputation to rely on long-term participation of established nodes. Such nodes help new nodes that join the system to overcome long-range attacks [13, 23], where an adversary presents an alternative history, indistinguishable from the true one. Additionally, in smaller systems, the value of tokens and hence the collateral amounts are volatile, leaving the system vulnerable to hijacking by a well-funded adversary.
To overcome this, stake can be managed on the primary blockchain, taking advantage of its decentralization and reliability: The nodes maintaining the expansion blockchain first lock their stake on the primary chain. Then, they agree on the series of blocks using a consensus protocol [16]. Managing stake for external tasks is already possible on major blockchains. EigenLayer [60] on Ethereum and Subnets [38] on Avalanche [52] are prominent examples. If nodes misbehave, they are penalized by losing their deposit on the primary chain. The underlying consensus protocol of the primary chain guarantees safety and liveness assuming that a sufficient majority of nodes are correct, i.e., follow the protocol.
Prior work (§2) extensively explored reconfiguration of nodes in Byzantine Fault Tolerant (BFT) systems [44]. Expansion chains like Polygon [42] and Avalanche subnets [38] have adopted these approaches. However, these systems rely on strong assumptions, such as synchronous communication among nodes or that nodes remain correct indefinitely. In practice, these assumptions may not hold: A user can unstake on the primary chain at any time, removing their collateral and hence their incentive to follow the expansion chain protocol. Thus, a node that is active when it is defined in a primary block might not be active by the time an expansion chain block is generated. Additionally, like the aforementioned PoS protocols, such expansion chains remain vulnerable to long-range attacks.
To address these challenges, particularly the vulnerability to long-range attacks and the potential for nodes to become Byzantine after unstaking, we first outline a model (§3). Nodes stake and unstake using the primary chain. Unstaking does not immediately withdraw the stake, but only after a delay $\Delta_{\textit{active}}$. A set of nodes that have stake locked at the same time are called a committee. Until any of them completed unstaking, the committee is active. As is standard [34, 50], we assume that at all times a sufficient majority of the nodes in each active committee are correct. The model employs a hybrid communication approach: nodes communicate asynchronously with each other until an unknown time $t_{\textit{GST}}$, after which message propagation is bounded, as in the partial synchrony model [25]. Nevertheless, at all times, nodes have synchronous access to read the current state of the primary chain (as when running an Ethereum node) and can write to it in bounded time (cf. [33]). This combination leverages the reliability of established blockchains while accounting for potential delays in internode communication in new protocols.
We present Aegis [^1] Named after the powerful divine shield carried by Zeus and Athena in Greek mythology, which first appears in Homer’s Iliad. (§4), an expansion blockchain that derives its security from a primary chain. It could be used to provide additional functionality to a primary chain, or more generally, for a blockchain unrelated to the primary chain, other than using its tokens as collateral.
The protocol includes a primary-chain contract that allows nodes to checkpoint the Aegis state. The nodes construct a blockchain in which each Aegis block references both its predecessor and a corresponding primary block (Figure 1). The committee defined by that primary block is the committee in charge of generating the next Aegis block using a consensus protocol. While maintaining the Aegis chain, its nodes occasionally checkpoint the latest block with the primary-chain contract. In the happy flow, the checkpoints are less than $\Delta_{\textit{active}}$ apart, so there is always an active committee.
@@
Figure 1: Aegis and primary-chain blocks.
@@
However, creating an Aegis block might take too long due to internode communication delays. If it takes longer than $\Delta_{\textit{active}}$, there would be nothing to checkpoint. In this case, any Aegis node can issue a reset on the primary chain to specify a new committee. At this point, the committee is defined in the primary-chain reset, but not yet in any Aegis block, until one is generated. The risk is that the committee defined by the reset and the one defined by the previous block both generate a block at the same height (distance from chain root), forming a so-called fork. To prevent that, both the protocol and the primary-chain contract enforce timing constraints, avoiding consensus decisions that cannot be checkpointed and premature resets.
@@
Aegis can be readily implemented to decentralize various services as expansion chains. We conclude (§6) with practical considerations for performance and security.
2 Related Work
To the best of our knowledge, previous work on dynamic committees and reconfiguration in BFT systems did not consider nodes that stop being correct after their tenure (§2.1). Various work on expansion chains addressed the problem under this assumption (§2.2) and overcoming long-range attacks with checkpointing (§2.3). Others utilize multiple chains to form a new one (§2.4). Aegis can be used by Layer-2 rollups that aggregate transactions (§2.5), but it solves a distinct problem.
2.1 Classical Reconfiguration
The reconfiguration problem received intensive treatment in the classical distributed-systems literature [2]. In this setting, committees are not defined by stake on the primary chain, as in our setting, but are rather defined by an external process.
Lamport et al. [43] implement reconfiguration in a crash-fault model using a separate, trusted configuration master. The nodes in one committee can trigger a reconfiguration by calling the configuration master, allowing the protocol to handle many failures by pruning out crashed nodes. However, this approach relies on having the new committee obtain the previous state from the previous committee, which is not reliable in a Byzantine model, particularly in our model where after deprecation the old committee behaves arbitrarily.
Aubin et al. [4] use black-box individual consensus algorithms. On reconfiguration, they require the previous committee to agree on which entries are aborted, such that this committee will not decide on them. This is also not possible in our model, where the nodes might become deprecated due to unstaking before completing the abort.
Abraham and Malkhi [2] identify that reconfiguring protocols rely on wedging, where each consensus result is stored by a non-byzantine replica. This replica can be, e.g., a larger set of replicas; it can be mapped to the primary chain in our model. They propose a synchronous reconfiguration protocol, but it relies on bounded latency among nodes to identify the wedging state. This is not possible in our model, where due to asynchrony previous committees might be deprecated before taking part in this protocol.
PBFT [16] and Zyzzyva [35] work in a fully asynchronous environment and use ${3f+1}$ replicas for reconfiguration. However, their progress relies on access to this reconfiguration master, implying that in case of reconfiguration events, the latency is at least long enough to read and write to the primary chain, even after $t_{\textit{GST}}$. This is not acceptable for our purposes. On the other hand, note our model is stronger: We assume synchronous access to the primary chain.
Matchmaker [62] assumes crash-fault only, and that nodes can access deprecated nodes. They explicitly assume that joining nodes access a matchmaker (implemented as a set of processes, mapped to our primary chain) to obtain the current state. All access is asynchronous, allowing to complete the access to the matchmakers. Mapping to our model, this would imply that the time to create every block is at least the time to read and write to the primary chain. Again, this is unacceptable for our purposes.
Kuznetsov and Tonkikh [36] consider slow clients, that might access a deprecated committee that became Byzantine and can behave arbitrarily. To overcome this, nodes are required to destroy deprecated private keys. However, this assumption does not apply to our model. We can’t assume that nodes would delete their old keys, since in our model, we assume that nodes of an active committee only follow the protocol due to the threat of slashing. But, slashing for not deleting old keys is not possible, as no one can tell when a node doesn’t delete its old keys.
Note that the cooldown period for unstaking, $\Delta_{\textit{active}}$, does allow bounded-time access to previous committees that are still active. This assumption is weaker than in prior work where previous committees remain active indefinitely. Nevertheless, it allows us to achieve the desired guarantees with Aegis.
2.2 Expansion Chains
Expansion chains are blockchains that rely on a primary chain for security. They can be implemented using EigenLayer [60], Avalanche subnets [38] and Polygon 2.0 [42]. Compared to a new blockchain, an expansion chain is easier to bootstrap, as it can rely on the security of the primary chain. Nodes can be required to stake on the primary chain, using its tokens as collateral to ensure their good behavior. Due to the established security of the primary chain, the value of its tokens is more stable, and the nodes are more likely to remain correct.
On the surface, securing an expansion chain seems like the classical State Machine Replication (SMR) [45] with dynamic committee reconfiguration [44]. Essentially, the committee for expansion block $k$ agrees not only on the block but also on the committee for the subsequent expansion block $k+1$, as defined in a primary block. However, this approach relies on an assumption that does not hold in our case, namely, that a correct node remains correct indefinitely. But the generation of expansion block $k+1$ could take arbitrarily long; meanwhile, the members of its committee might unstake in the primary chain, so they are no longer incentivized by collateral. The result is either a deadlock (if we ignore blocks by a deprecated committee) or a block generated by a committee that is no longer incentivized to follow the protocol. Aegis addresses this issue by using references from the expansion blocks to the primary blocks to define the committee for the next block, and allowing for a reset if the committee becomes obsolete before it generates a block.
BMS [57] is a mechanism for secure reconfiguration of blockchains via a trusted platform that manages reconfiguration process. However, it requires expansion chains to approve each reconfiguration, which is not always possible: If a node wishes to withdraw their stake, it is not generally acceptable to wait for such an approval, particularly if the expansion chain is asynchronous and the time is unbounded. Aegis does not require such approval, and nodes can independently withdraw without affects its security.
Avalanche subnets [38] are blockchains whose validators are defined, added and removed by transactions on a primary chain. Avalanche proposes this mechanism and facilitates open participation in its so-called elastic subnets. EigenLayer [60] allows Ethereum validators to restake their Ethereum-protocol deposits, committing to validating a secondary system with the same stake used for Ethereum, meaning that the same deposit can be slashed due to misbehavior in either system. Both mechanisms provide a new tool for creating expansion protocols, but leave open the question of securing the protocol, which Aegis addresses as mentioned above.
Polygon 2.0 [42] secures a secondary blockchain using Ethereum as primary blockchain. It uses staking on Ethereum to define the nodes maintaining separate blockchains. However, it encounters the aforementioned issue of SMRs with dynamic reconfiguration, meaning its security implicitly relies on either synchronous communication among Polygon nodes or that correct nodes remain correct indefinitely. This is because node set changes are events on Ethereum [40] and are voted in Polygon [41]. Therefore, if it can take arbitrarily long to reach consensus and members are not indefinitely correct, by the time they reach consensus they might not be correct. In contrast, Aegis does not rely on either assumption, while always ensuring safety and ensuring progress after $t_{\textit{GST}}$.
Merge-mining is another approach to secure expansion blockchains [11]. It is a technique for Proof-of-Work (PoW) systems [46, 55], where nodes are called miners and they generate blocks by guessing hash preimages. In merge-mining, a Proof of Work miner mines on two PoW blockchains with the same computational power: each hash it calculates can result in a block for either chain. This allows the computational power securing a major PoW chain to also secure an expansion chain. However, if a miner misbehaves when generating an expansion chain block, it suffers no penalty on the main chain, which remains oblivious. In Aegis such misbehavior is penalized, so nodes are strongly incentivized to follow the protocol.
2.3 Long-Range Attacks and Checkpoints
Long-range attacks are a type of attack where an adversary presents an alternative history, indistinguishable from the true one, to the network. This can be used to double-spend coins, or to prevent the network from reaching consensus. Newcomers to the network are vulnerable to these attacks, as they do not have the full history of the blockchain.
Long-range attacks are possible in both PoW and PoS blockchains. In PoW, an adversary performing a long-range attack is required to hold a majority of the computational power, and use it to mine a longer chain than the honest chain [^2] More specifically, this is called a 51% attack [8].. However, this form of attack is more severe in PoS, as it doesn’t require significant computational power, only a majority of the stake at a given time. This is exacerbated by the fact that old nodes, who have already withdrawn their stake, are no longer at risk of being slashed.
Previous work has explored checkpoints to prevent long-range attacks. Every several blocks, a block is checkpointed by nodes, in order for newcomers to later retrieve the latest checkpoint to be able to distinguish between the current chain and a long-range attack. Earlier work considered social consensus [9, 13, 21], where a committee of participants, independent of the blockchain, agrees on the current state of the blockchain and communicates it to others via some external medium. Later, some works utilized a BFT consensus protocol to perform the checkpointing [51, 53]. Afterward, the idea of using a separate proof-of-work blockchain to timestamp checkpoints was introduced [32, 6, 59]. These works suggested injecting a digest of the checkpoint into a secure PoW blockchain as a transaction, thus timestamping the checkpoint. Another related work is Winkle [5], which suggests weighing checkpoints based on the volume of transactions which were place in blocks following the checkpoint.
Similarly, Aegis enjoys protection from long-range attacks due to its checkpoints. But unlike them, even if the expansion chain is stalled, reconfiguration occurs in the primary chain, and Aegis ensures that this will be reflected in the expansion chain. None of the aforementioned solutions can guarantee this property, as they all implement reconfiguration within the expansion chain.
2.4 Chain Combining

### v3

# Aegis: Tethering a Blockchain with Primary-Chain Stake

@@
Abstract.
Blockchains implement decentralized monetary systems and applications. Recent advancements enable what we call tethering a blockchain to a primary blockchain, securing the tethered chain by nodes that post primary-chain tokens as collateral. The collateral ensures nodes behave as intended, until they withdraw it. Unlike a Proof of Stake blockchain which uses its own token as collateral, using primary-chain tokens shields the tethered chain from the volatility of its own token.
State-of-the-art tethered blockchains either rely on centralization, or make extreme assumptions: that all communication is synchronous, that operators remain correct even post-withdrawal, or that withdrawals can be indefinitely delayed by tethered-chain failures.
@@ Abstract.
We prove that with partial synchrony, there is no solution to the problem. However, under the standard assumptions that communication with the primary chain is synchronous and communication among the tethered chain nodes is partially synchronous, there is a solution. We present a tethered-chain protocol called Aegis. Aegis uses references from its blocks to primary blocks to define committees, checkpoints on the primary chain to perpetuate decisions, and resets to establish new committees when previous ones become obsolete. It ensures safety at all times and rapid progress when latency among Aegis nodes is low.
Byzantine Consensus, BFT, Byzantine fault tolerance, Blockchain, Reconfiguration
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 13–17, 2025; Taipei, Taiwanisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Distributed computing methodologiesccs: Security and privacy Distributed systems security
1. Introduction
@@
Blockchains, such as Ethereum (Wood, 2015), implement monetary systems and allow users to deploy smart contracts (Szabo, 1996) with arbitrary logic. Users issue transactions to exchange assets and interact with smart contracts. Blockchain nodes aggregate transactions into blocks, forming a chain that determines the system’s state.
As blockchain adoption grows, scalability has emerged as a critical challenge (Chauhan et al., 2018; Zhou et al., 2020). Blockchains have limited throughput, often high latency, and are restricted to deterministic execution. To address these limitations and implement new functionality, new blockchains are created (coi, 2024). These new chains are either independent or Layer-twos (L2s) that expand existing blockchains [^1] Not all L2s are blockchains by themselves, but many are (L2BEAT, 2025)..
A popular approach to secure blockchains is Proof of Stake (PoS). In a PoS blockchain, nodes can stake, namely, lock their tokens as collateral (Kiayias et al., 2017; Daian et al., 2019; D’Amato et al., 2022). Until they unstake, namely, withdraw their collateral, they are active nodes, participating in the creation of new blocks. If they misbehave, they are penalized via slashing (Buterin, 2015), losing all or part of their collateral.
PoS works well for major blockchains that have reached a critical mass of participants and assets (coi, 2024). Yet in smaller nascent systems, the token value is not established and depends on the blockchain’s success, which in turn requires the blockchain to be secure. However, a PoS blockchain can only be considered secure if the value of its collateral, expressed in its tokens, is high enough. This chicken-and-egg problem poses an obstacle to newer chains, and remains a concern afterward, since token prices are often volatile (coi, 2024).
New systems like EigenLayer (EigenLayer, 2023) and Symbiotic (Symbiotic, 2025) on Ethereum and Babylon (Team, [n. d.]) on Bitcoin (Bitcoin community, 2013) enable staking tokens for securing external tasks. Some of these, such as EigenLayer, implement restaking (Durvasula and Roughgarden, 2024), allowing nodes to simultaneously secure both their native blockchain and additional tasks, while others, like Babylon, dedicate stake exclusively to new tasks. These stake management systems represent a promising direction and have already garnered assets worth billions of US dollars (DefiLlama, 2025).
We refer to tethering as using a primary blockchain for defining and coordinating a secondary chain’s committee, a set of nodes that create new blocks in the tethered chain using a consensus protocol (Castro et al., 1999). Stake management systems enable a particularly useful form of tethering, where the primary chain’s tokens serve as collateral, shielding the tethered chain from the volatility of its own token. Beyond addressing token volatility, tethering provides a practical solution for secure committee changes (epoch transitions), a notorious challenge in blockchain design (Bano et al., 2019). Throughout this paper, we use stake-based terminology to discuss tethering, where nodes stake tokens to join committees and unstake to exit. Like PoS blockchains, nodes are incentivized to be correct, i.e., follow the protocol, under the threat of slashing, and may unstake at any time.
@@
[…]
Illustration of Aegis and primary-chain blocks.
Figure 1. Aegis and primary-chain blocks.
Prior work (§2) extensively explored reconfiguration of nodes in Byzantine Fault Tolerant (BFT) systems (Lamport et al., 2010). A common approach is for an old committee to select the new committee before its tenure ends. However, in our setting, the old committee might become deprecated before it can select the new committee: When a node unstakes on the primary chain, this should trigger a reconfiguration, but if the committee fails to reach consensus in time, the node becomes unstaked and thus no longer incentivized to be correct.
State-of-the-art tethered chains, including recent work by Dong et al. (Dong et al., 2024) and operational systems like Polygon (Labs, 2023b), Avalanche Subnets (Labs, 2025) and Cosmos (Cosmos, 2025), have adopted the classical approach. These systems either disallow unstaking until a successful reconfiguration completes—risking validators’ collateral being locked indefinitely if reconfiguration fails—or rely on strong assumptions to avoid the failure scenario described above: They either assume synchronous message delivery with bounded time, ensuring committees reach consensus in time, or assume nodes remain correct indefinitely even after unstaking, making further reconfiguration unnecessary.
Various Layer-2 solutions also aim to extend blockchain functionality and performance (Gudgeon et al., 2020). Rollups batch transactions and post summaries to the primary chain (Kalodner et al., 2018; Arbitrum, 2023; Optimism, 2023; Starkware, 2023; Labs, 2023a), but rely on centralized sequencers and are bound by primary-chain finality times. While recent work attempts to decentralize sequencers (Motepalli et al., 2023; Capretto et al., 2024), it doesn’t support dynamic committees. Other approaches include sidechains (Back et al., 2014) that operate as separate blockchains connected to another blockchain through bridges for asset transfer, sharding (Kokoris-Kogias et al., 2018; Wang et al., 2019) that modifies the primary chain’s consensus to partition validators, and payment channels (Poon and Dryja, 2015; Decker and Wattenhofer, 2015) that enable rapid off-chain bilateral transactions. Unlike these approaches, tethered chains aim to be independent decentralized chains with their own rapid progress rate that derive security from primary-chain stake, without requiring modifications to the primary chain.
To address the challenge of building a secure tethered chain, we first outline a model (§3). Nodes can stake and unstake using the primary chain, but unstaking only completes after a delay $\Delta_{\textit{active}}$. Nodes with locked stake at the same time form a committee, which remains active until any member completes unstaking. As is standard (Kiayias et al., 2017; Pu et al., 2023), we assume that at all times a sufficient majority of nodes in each active committee are correct. The model employs a hybrid communication approach: Nodes communicate asynchronously with each other until an unknown global stabilization time $t_{\textit{GST}}$, after which message propagation is bounded, as in the partial synchrony model (Dwork et al., 1988). However, at all times, nodes have synchronous (though slow) access to read and write to the primary chain (cf. (Kelkar et al., 2020)).
The choice of a hybrid communication model is not coincidental, but necessary to build a secure tethered chain (§4). To account for real-world conditions, where there may be potential outages in internode communication, especially in nascent systems, we assume partial synchrony. However, we prove that if communication with the primary chain is also partially synchronous, no solution can be secure. To circumvent this impossibility, we bound the access-time to the primary chain, similarly to many existing L2s (Optimism, 2023; Arbitrum, 2023). Nevertheless, this bound can be chosen to be sufficiently large (possibly in the order of hours or days), and should not affect the progress rate of the tethered chain.
We present Aegis [^2] Named after the powerful divine shield carried by Zeus and Athena in Greek mythology, which first appears in Homer’s Iliad. (§5), a tethered blockchain. Each Aegis block references both its predecessor and a corresponding primary block. The committee defined by that primary block is the committee in charge of generating the next Aegis block using a consensus protocol. While maintaining the Aegis chain, its nodes occasionally checkpoint the latest block with a dedicated smart contract on the primary-chain.
@@
In the happy flow, the checkpoints are less than $\Delta_{\textit{active}}$ apart, so there is always an active committee. However, creating an Aegis block might take too long due to internode communication delays. If it takes longer than $\Delta_{\textit{active}}$, there would be nothing to checkpoint. In this case, any Aegis node can issue a reset on the primary chain to specify a new committee. At this point, the committee is defined in the primary-chain reset, but not yet in any Aegis block, until one is generated.
There is a risk that the committee defined by the reset and the one defined by the previous block both generate a block at the same height (distance from chain root), forming a so-called fork. To prevent that, both the protocol and the primary-chain contract enforce timing constraints, avoiding consensus decisions that cannot be checkpointed and premature resets.
The Aegis protocol thus comprises the primary-chain smart contract, which handles checkpoints and resets, and a distributed protocol executed by Aegis nodes, which invokes a classical single-committee consensus protocol to extend the Aegis chain.
@@
We then discuss practical considerations for performance and security, as well as potential extensions of our tethering paradigm (§7). While our implementation focuses on stake-based committee management, the reset mechanism can be adapted to support different committee transition logics, such as majority decisions or after a predetermined number of blocks.
Aegis can be readily implemented to deploy various decentralized services as tethered chains. These include scaling solutions, like rollups; and solutions providing new functionality, like oracles and randomness sources.
2. Related Work
Classical Reconfiguration.
The reconfiguration problem has received intensive treatment in the classical distributed-systems literature. Some solutions rely on the old committee completing a full state handoff before leaving (Lamport et al., 2009), or finalizing and aborting any pending entries (Aublin et al., 2015); both assumptions fail if nodes exit prematurely. Other approaches store consensus decisions in a separate reliable replica (Abraham and Malkhi, 2016; Castro et al., 1999; Kotla et al., 2007; Whittaker et al., 2021), incurring frequent reads and writes to the replica (Castro et al., 1999; Kotla et al., 2007; Whittaker et al., 2021), which implies additional latency. In our setting, a cooldown period for unstaking provides a bounded window to contact the outgoing committee via the primary chain, eliminating the need for continuous access to a reconfiguration service. Another proposal requires deprecated committees to destroy old private keys (Kuznetsov and Tonkikh, 2022), but this can’t be enforced in a stake-based model.
Previous work on state machine replication we are aware of sufficed with external validity. That is, orders that are correctly formatted are considered valid, and the only requirement is that such orders are eventually placed in the log. This allows values to be aggregated as well as submitted long before they are placed in the blockchain (Castro et al., 1999). This assumption is insufficient for blockchain applications, where various, possibly competing users submit transactions whose content depends on the latest system state. We therefore introduce a stronger validity requirement called no-batching validity.
Tethered Chains.
Tethered chains are blockchains that use tokens from a primary chain as collateral, simplifying bootstrapping compared to starting a new blockchain and shielding them from the volatility of their own token. At first glance, tethered chains may resemble classical State Machine Replication with dynamic reconfiguration (Lynch, 1996; Lamport et al., 2010), where the committee for block $k$ selects the committee for block $k+1$. However, if the creation of block $k+1$ is delayed, members of its committee may unstake in the meanwhile and lose their incentive to behave correctly. Aegis addresses this by having each block reference its next committee directly from the primary chain and allowing a reset if the current committee stalls.
Existing tethered systems, including Avalanche subnets (Labs, 2025), Polygon 2.0 (Labs, 2023b), and Cosmos (Cosmos, 2025), face this fundamental issue. They either assume synchronous communication to avoid delayed block creation, or assume nodes remain correct indefinitely so delayed blocks are still handled correctly. New tethered chains can be created using stake management solutions such as EigenLayer (EigenLayer, 2023) and Symbiotic (Symbiotic, 2025) on Ethereum or Babylon (Team, [n. d.]) on Bitcoin, these platforms leave open how to implement the chain.
BMS (Steinhoff et al., 2021) proposes a trusted reconfiguration platform that must approve each committee change, which becomes impractical if a node wants to exit but the tethered chain fails to grant timely approval. By contrast, Aegis enables nodes to withdraw unilaterally without compromising security, thanks to primary-chain collateral tracking.
An alternative approach, merge-mining (Bonneau et al., 2015), applies to Proof-of-Work systems (Nakamoto, 2008; Sompolinsky et al., 2021) by reusing computational power across blockchains, but provides no penalty on the main chain for misbehavior on the tethered chain. Because Aegis slashes misbehaving nodes at the primary-chain level, it preserves correct incentives even after committees become deprecated.
Long-Range Attacks and Checkpoints.
Long-range attacks occur when an adversary presents an alternative history that appears indistinguishable from the genuine chain, particularly to newcomers lacking the full chain state (Buterin, 2014a). These attacks enable double-spending or can disrupt consensus. They can affect both PoW and PoS blockchains, though in PoW an attacker must hold a majority of the network’s computational power (a 51% attack (Bano et al., 2019)). In PoS, however, only a majority of stake at some historical point is required, and once nodes withdraw, they are no longer at risk of slashing. As a result, long-range attacks are often considered more severe in PoS settings.
Prior work has explored checkpointing to mitigate such attacks. Early methods relied on social consensus, wherein an external committee vouches for the current chain state (Barber et al., 2012; Buterin, 2014b; Daian et al., 2019). Later studies employed BFT protocols for checkpoint creation (Rana et al., 2022; Sankagiri et al., 2021), or used separate PoW networks to timestamp checkpoints (Karakostas and Kiayias, 2021; Azouvi and Vukolić, 2022; Tas et al., 2023). Because Aegis already depends on the primary chain for stake management, it reuses that same chain to record checkpoints, avoiding extra trust assumptions. Another approach, Winkle (Azouvi et al., 2020), suggests weighting checkpoints by the volume of subsequent transactions.
Layer-2 Solutions.
Various Layer-2 (L2) solutions aim to extend the performance and functionality of a primary chain (also called Layer-1 or L1) but differ fundamentally in their approach and limitations. Rollups improve transaction throughput by batching transactions in a separate ledger and occasionally posting summaries to the L1 for committing changes. Rollups come in two forms: Optimistic rollups (Kalodner et al., 2018; Arbitrum, 2023; Optimism, 2023) rely on a dispute process in the L1 to correct, and possibly penalize, incorrect summaries; while zero-knowledge rollups (Starkware, 2023; Labs, 2023a) generate cryptographic proofs of summary validity. Gudgeon et al. (Gudgeon et al., 2020) provide a comprehensive systemization of knowledge.
A key limitation of rollups is their reliance on centralized components for transaction ordering (via sequencers), which poses a risk of transaction censorship and manipulation for financial gain. Recent work has explored decentralizing the sequencer (Motepalli et al., 2023), with Setchain (Capretto et al., 2024) implementing Byzantine-tolerant transaction ordering, but its use of static committees limits its ability to handle dynamic stake changes. Another limitation is that transaction finality is bound to block confirmation times in the primary chain, as changes are only finalized once a summary is finalized in the primary chain. Aegis overcomes both limitations with its decentralized consensus and rapid finality: Blocks can be accepted once they are created, which may be significantly sooner than until they are checkpointed. Furthermore, Aegis can serve as a decentralized sequencer for rollups with dynamic committee reconfiguration.
Sidechains (Back et al., 2014; Polygon, 2023; Gnosis, 2023) operate as independent blockchains that connect to another blockchain through bridges, protocols that enable asset transfers between chains. These are orthogonal to tethered chains, as a tethered chain can also be a sidechain if it allows asset transfer.
Sharding (Kokoris-Kogias et al., 2018; Wang et al., 2019) takes a more invasive approach by modifying the core consensus protocol of the L1 to partition validators across multiple chains. Unlike Aegis, which operates independently while deriving security from the primary chain, sharding requires fundamental changes to the L1 design.
Payment channels are point-to-point channels (Poon and Dryja, 2015; Decker and Wattenhofer, 2015) that can aggregate several payments between two parties and form payment networks. While they enable efficient transactions between parties, they fundamentally differ from Aegis’s approach as they rely on bilateral relationships rather than decentralized consensus, with each channel exclusively controlled by its two counterparties.
Our goal differs from these approaches: We seek to create a chain with its own rapid progress rate, which relies on a primary chain for managing its stake; other than this dependency, it may not have any other relation to the primary chain.
Chain Combining.